Find an equation of the line containing the given pair of points ((-1,-3) and (-9,-2). The equation of the line in slope intercept form is y=
Answer by stanbon(75887)

Find an equation of the line containing the given pair of points
((-1,-3) and (-9,-2). The equation of the line in slope intercept form is y=
--------
slope = (-2--3)/(-9--1) = 1/-8 = -1/8
-------------------
Form y = mx + b
----
-3 = (-1/8)(-1) + b
b = -24/8 - (1/8)
b = -25/8
Equation:
y = (-1/8)x - (25/8)
============================
Cheers,
Stan H.